### v1

# View-aware Cross-modal Distillation for Multi-view Action Recognition

Abstract
The widespread use of multi-sensor systems has increased research in multi-view action recognition. While existing approaches in multi-view setups with fully overlapping sensors benefit from consistent view coverage, partially overlapping settings where actions are visible in only a subset of views remain underexplored. This challenge becomes more severe in real-world scenarios, as many systems provide only limited input modalities and rely on sequence-level annotations instead of dense frame-level labels. In this study, we propose View-aware Cross-modal Knowledge Distillation (ViCoKD), a framework that distills knowledge from a fully supervised multi-modal teacher to a modality- and annotation-limited student. ViCoKD employs a cross-modal adapter with cross-modal attention, allowing the student to exploit multi-modal correlations while operating with incomplete modalities. Moreover, we propose a View-aware Consistency module to address view misalignment, where the same action may appear differently or only partially across viewpoints. It enforces prediction alignment when the action is co-visible across views, guided by human-detection masks and confidence-weighted Jensen–Shannon divergence between their predicted class distributions. Experiments on the real-world MultiSensor-Home dataset show that ViCoKD consistently outperforms competitive distillation methods across multiple backbones and environments, delivering significant gains and surpassing the teacher model under limited conditions.
1 Introduction
The increasing use of multi-sensor systems in smart homes [nguyen2025multisensor], surveillance [li2022attention], and assistive robotics [abdelaal2020multi] has driven growing interest in multi-view action recognition. By aggregating observations from multiple viewpoints, such systems capture richer spatio-temporal cues than single-view setups and improve robustness to occlusion, viewpoint changes, and background clutter [wang2014cross, shahroudy2016ntu, liu2019ntu, olagoke2020literature]. Most existing methods, however, are designed for fully overlapping sensor setups. In such settings, sensors cover the same region of the scene, and each action instance is fully visible across all views. Datasets such as Nanyang Technological University (NTU) RGB+D [shahroudy2016ntu, liu2019ntu] and NorthWestern-University of California at Los Angeles (NW-UCLA) Multiview Action 3D [wang2014cross] adopt this assumption, enabling methods [das2023viewclr, liu2023dual, shah2023multi, yang2026trunk] to exploit contrastive or disentanglement learning to enforce viewpoint invariance. In contrast, partially overlapping settings are more representative of real-world deployments. These setups feature non-aligned sensor coverage, where an action may be visible in some views but completely occluded or out of scope in others. Recent methods [yasuda2022multi, yasuda2024guided, nguyen2024action, nguyen2025multisensor] address this challenge by proposing fusion strategies to integrate fragmented evidences or recover missing observations. However, they focus primarily on feature and view fusion, without explicitly enforcing consistency across views when actions are only partially observed. Therefore, their effectiveness in realistic partially overlapping scenarios is limited, and explicit modeling of view-aware consistency remains underexplored.
On the other hand, multi-view action recognition models achieve their best performance when all modalities are fully available and dense frame-level annotations are provided [nguyen2025multisensor]. However, in real-world multi-sensor deployments, these conditions are often unmet. First, modality-limited scenarios occur when some input streams (e.g., audio or depth) are unavailable due to hardware or environmental constraints [10.1145/3711860]. Second, annotation-limited scenarios arise as frame-level labeling is prohibitively expensive for multi-view videos, leaving only weak sequence-level tags in many datasets [yasuda2022multi, yasuda2024guided]. Although prior works [john2024frame, nguyen2024action, 10.1145/3744742] address learning under limited conditions, they primarily rely on weak supervision, remaining sensitive to fragmented observations and noisy labels.
To handle this limitation, the potential direction is to transfer supervision from a stronger model to a weaker one. Knowledge Distillation (KD) [hinton2015distilling] offers a promising solution by allowing a fully supervised multi-modal teacher trained with strong labels to guide a modality- and annotation-limited student. Multi-modal KD (MKD) [gupta2016cross, garcia2021distillation, chen2021distilling, perez2020audio] extends this concept to cases where the teacher benefits from multi-modal inputs and frame-level labels, while the student operates with incomplete modalities or only coarse sequence-level labels. However, applying KD to partially overlapping multi-view action recognition remains challenging, as it requires transferring knowledge while preserving consistency across views where the action may appear in some views but be absent in others.
To address these challenges, we propose View-aware Cross-modal Knowledge Distillation (ViCoKD), a method that distills knowledge from a fully supervised multi-modal teacher to a modality- and annotation-limited student. ViCoKD employs feature-level and logit-level KD, leveraging a cross-modal adapter that enables the student to exploit audio-visual correlations through Cross-modal Attention while operating solely on visual input. In addition, we introduce a View-aware Consistency module that explicitly handles partially overlapping sensor setups by encouraging the model to produce the same prediction across different views whenever the action is co-visible. This is achieved through a confidence-weighted Jensen–Shannon divergence [menendez1997jensen] between the predicted class distributions of co-visible views, ensuring that supervision is concentrated on frames with reliable, view-consistent evidence. The main contributions of this study are as follows:
ViCoKD method: A cross-modal attention-based KD method that transfers knowledge from a fully supervised multi-modal teacher to a modality- and annotation-limited student for multi-view action recognition in partially overlapping sensor settings.
View-aware Consistency module: A consistency module that leverages human-detection masks and confidence-weighted Jensen–Shannon divergence [menendez1997jensen] to align predictions only on view pairs with reliable and visible action evidence.
Comprehensive evaluation: Experiments on the real-world MultiSensor-Home dataset [nguyen2025multisensor] across diverse backbones and home environments show that ViCoKD consistently outperforms competitive distillation methods and surpasses the teacher under limited conditions.
2 Related Work
Multi-view Action Recognition. Depending on sensor coverage, multi-view action recognition tasks are studied under either fully overlapping or partially overlapping settings. Most prior work focuses on the former, with datasets such as NTU RGB+D [shahroudy2016ntu, liu2019ntu], NW-UCLA [wang2014cross], and Toyota Smarthome [das2019toyota] where all views capture the same subject. Building on these datasets, ViewCLR [das2023viewclr] employs view generation to generalize to unseen viewpoints, ViewCon [shah2023multi] applies supervised contrastive learning to produce viewpoint-invariant embeddings, and Dual-Recommendation Disentanglement Network (DRDN) [liu2023dual] leverages disentanglement learning to decouple action-specific and view-specific cues. In contrast, real-world applications involve wide-area surveillance with only partially overlapping views, where actions may be visible from some sensors but entirely occluded in others. Recent work has introduced datasets tailored to this challenging setting [yasuda2024guided, nguyen2025multisensor]. Yasuda et al. [yasuda2022multi] introduce MultiTrans, which models inter-sensor relationships to integrate multi-view cues, and later propose Guided Masked sELf-Distillation (Guided-MELD) [yasuda2024guided] to handle fragmented observations by reducing redundancy and filling missing sensor data for coherent event-level representations. John and Kawanishi [john2024frame] present a weakly supervised latent embedding framework that learns from sequence-level labels while enabling frame-level action detection. Recently, Nguyen et al. [nguyen2024action, nguyen2025multisensor] proposed MultiASL for weakly supervised action selection across sensors and MultiTSF for robust cross-sensor temporal fusion. Despite these advances, most existing work focuses on feature and view fusion strategies, while explicit modeling of view-aware consistency in partially overlapping settings is still limited.
Multi-modal Knowledge Distillation (MKD). KD [hinton2015distilling] trains a smaller or weaker student model to mimic the predictions or intermediate representations of a teacher model, improving efficiency or performance in limited conditions. Beyond matching final logits [hinton2015distilling], KD has evolved to include feature-level supervision [romero2015fitnetshintsdeepnets, zagoruyko2016paying] and relational knowledge transfer [park2019relational] to strengthen teacher-student alignment. MKD [chen2021distilling, zhang2021knowledge] extends KD to settings where the teacher and student operate over multiple modalities. The teacher typically leverages full-modality inputs, such as RGB, audio, depth, or skeleton data, while the student learns to replicate its performance under restricted modalities. Prior works have demonstrated the benefits of MKD for RGB-depth [gupta2016cross], RGB-flow [garcia2021distillation], and audio-visual [chen2021distilling, perez2020audio] tasks. In video understanding, MKD has been shown to improve recognition under missing modalities [radevski2023multimodal] by transferring temporal and cross-modal cues from the teacher to the student. In this work, we study MKD in the context of partially overlapping multi-view action recognition under missing-audio conditions. Unlike these methods, our student operates solely with visual inputs and is trained under frame-level or sequence-level supervision, the latter representing a more challenging weakly supervised setting. To address this gap, we propose the ViCoKD method with View-aware Consistency, enabling students to learn robust cross-modal and cross-view representations under the limited conditions.
3 Proposed Method
We address the task of multi-view action recognition in partially overlapping settings within a KD framework. Actions are captured from synchronized viewpoints. In this study, we address the multi-modal setting as comprising audio and visual inputs. At each time step , the input from view of teacher comprises two modalities: audio segment , represented as a time-aligned spectrogram with frequency bins per frame, and visual frame , where and denote spatial resolution, and is the number of color channels. Audio and visual inputs from view are denoted as and , respectively. The objective is to train a student model for multi-label action classification by transferring knowledge from a fully supervised multi-modal teacher model. Here, we focus on the audio-missing setting, where the student lacks access to the audio modality. Two supervision scenarios are considered for the student: (1) frame-level supervision, where the student is trained with frame-level and sequence-level labels, and (2) sequence-level supervision, where the student is trained using only video-level labels. The goal is to exploit the feature representations learned by the teacher (Figure 1(a)) and distill this knowledge into a student model (Figure 1(b)) under missing-modality conditions.
3.1 Teacher Model with Cross-modal Attention
Figure 1(a) illustrates the teacher model, which is designed to learn spatio-temporal representations from synchronized multi-modal multi-view inputs, trained with strong supervision at the frame and sequence levels.
Modality-specific Encoders. Each input sequence consists of audio data and visual data from view , which are processed by modality-specific encoders. Specifically, we use Audio Spectrogram Transformer (AST) [gong2021ast] as the audio encoder and Vision Transformer (ViT) [dosovitskiy2020image] as the visual encoder :
where and are temporally aligned audio and visual feature sequences, respectively. The AST and ViT encoders share parameters across all views to ensure consistent representation learning.
Cross-modal Attention Module. To integrate audio-visual information, we adopt a cross-modal attention mechanism [ye2019cross, wei2020multi] where audio features are used as keys and values and visual features serve as queries. Positional encodings are added to both modalities to retain temporal and spatial ordering before cross-modal attention is applied as:
The attended features are fused with the visual features via element-wise addition as:
This fusion strategy preserves the spatial-temporal structure of the visual representation while enhancing it with complementary audio information.
Temporal Encoding. The fused representation from each view is fed into a Transformer-based [vaswani2017attention] temporal encoder to capture sequential dependencies across time as:
where TemporalEncoder is a multi-layer Transformer that models self-attention across frames to learn long-range temporal patterns within the sequence. Output represents a temporally enriched feature sequence encoding spatio-temporal cues.
Supervision Objective. The teacher is optimized with frame-level and sequence-level classification losses. At the frame-level, we predict per-frame logits and compute a Two-way loss [kobayashi2023two] that combines sample-wise and class-wise objectives. At the sequence-level, we apply temporal average pooling followed by classification, using the same loss formulation. The overall training objective for the teacher is calculated as:
where and are the frame-level and sequence-level classification losses, respectively, and is a balancing hyperparameter.
3.2 Teacher-to-student View-aware Cross-modal Knowledge Distillation
Figure 1(b) illustrates the knowledge distillation pipeline, where the student adopts the same architecture as the teacher while operating with a missing audio modality. The student is trained to mimic the teacher’s knowledge through the supervision objective in Eq. (5), combined with a KD loss that transfers feature- and logit-level representations from the teacher. In addition, the student benefits from the View-aware Consistency loss, which enforces consistent predictions across partially overlapping views when the action is co-visible.
Cross-modal Adapter. When audio is missing, a cross-modal adapter is employed to synthesize pseudo-audio features from visual features , where is a lightweight feed-forward network with a non-linear activation. The pseudo-audio feature is then used in cross-modal attention as:
Feature Distillation (FD). We distill knowledge at two stages: the attended audio-visual features from the cross-modal attention, which transfer audio cues into the visual stream, and the fused representations , which integrate multi-modal context. Formally, we minimize the frame-wise Euclidean distance between the teacher’s and the student’s features as:
This alignment enhances the student’s ability to inherit the teacher’s cross-modal attention behavior and multi-modal feature integration.
Logit Distillation (LD). We apply logit-level distillation at the frame-level on the fused features from multiple views. Let and denote the temperature-scaled softmax outputs of the teacher and student, respectively. The loss is computed as the Kullback–Leibler divergence [kullback1951information] (KL) between these distributions as:
where is the temperature parameter.
View-aware Consistency Module. To enforce consistent predictions across different views, we introduce a view-aware consistency loss applied at the frame-level. Figure 1(b) shows the details: for each view and time step , we obtain probabilities . A binary visibility mask is generated using human detection [wang2024yolov10], where indicates that a human is visible in view at frame , suggesting a high potential for action occurrence. Given a view pair , we compute the Jensen–Shannon divergence [menendez1997jensen] (JS) between their prediction distributions:
where . The divergence is computed only on frames where , and is weighted by the product of the maximum class probabilities from both views to emphasize confident predictions. The View-aware Consistency loss is defined as:
where serves as a confidence weight, encouraging consistency when both views are individually certain in their predictions, and is the number of view pairs. This loss aligns predictions where the action is visible in both views and places greater emphasis on high-confidence predictions.
4 Experiment
4.1 Experimental Conditions
Data Preparation. We conduct the experiments on the publicly available MultiSensor-Home dataset [nguyen2025multisensor], which includes two home environments: Home-1 and Home-2, as shown in Figure 2. Following the official experimental setup, we use a 7:3 train/test split within each environment. To ensure a balanced class distribution, the iterative stratification strategy proposed in [sechidis2011stratification] is employed. For all experiments, we extract a fixed number of synchronized frames per sample. During training, we apply uniform sampling with slight random perturbations to generate frame sequences that cover the entire video while preserving the fixed length . This sampling strategy also augments temporal data, enhancing robustness by introducing sequence variability. For evaluation, we adopt deterministic uniform sampling without perturbation to ensure reproducibility across runs.
Evaluation Metrics. Following [kobayashi2023two, nguyen2024action, nguyen2025multisensor], we evaluate multi-label action recognition performance using mean Average Precision (mAP), a macro-averaged metric widely adopted in multi-label classification tasks.
MultiTrans [yasuda2022multi] | MultiASL [nguyen2024action] | MultiTSF [nguyen2025multisensor] | |||||
| AV V | AV V | AV V | AV V | AV V | AV V | ||||
| Home-1 | Teacher | 61.40 | 73.81 | 76.12 | |||||
| Student | 60.77 | 57.59 | 63.24 | 55.91 | 73.17 | 61.17 | |||
| LogitKD | 61.28 | 60.48 | 65.26 | 58.78 | 70.08 | 60.67 | |||
| ModalKD | 58.22 | 54.04 | 69.50 | 58.92 | 73.67 | 62.88 | |||
| DualKD | 58.84 | 55.57 | 68.42 | 60.45 | 73.12 | 63.22 | |||
| CoKD | 61.34 | 60.65 | 75.33 | 64.52 | 76.89 | 63.27 | |||
| ViCoKD (Ours) | 66.27 | 62.66 | 76.83 | 65.23 | 82.91 | 65.28 | |||
| Home-2 | Teacher | 86.60 | 90.14 | 92.12 | |||||
| Student | 84.06 | 80.83 | 82.25 | 81.67 | 86.89 | 79.16 | |||
| LogitKD | 86.70 | 85.55 | 87.59 | 84.94 | 88.60 | 84.73 | |||
| ModalKD | 85.43 | 83.35 | 91.45 | 87.98 | 90.92 | 87.56 | |||
| DualKD | 86.41 | 84.63 | 85.58 | 81.88 | 87.71 | 85.36 | |||
| CoKD | 88.64 | 86.12 | 91.77 | 86.97 | 88.42 | 87.75 | |||
| ViCoKD (Ours) | 90.86 | 89.31 | 91.80 | 89.89 | 91.27 | 89.21 | |||
Comparison Methods. To evaluate the effectiveness of the proposed ViCoKD method, we compare against several representative KD methods under a consistent experimental setup. All methods adopt the same teacher–student configuration for a fair comparison. The teacher network is trained with frame-level audio-visual modalities (AVF), while the student is trained with visual-only inputs under either frame-level or sequence-level supervision (VF or VS). The teacher and student models use the architectures from the original MultiTrans [yasuda2022multi], MultiASL [nguyen2024action], and MultiTSF [nguyen2025multisensor] methods.
Student/Teacher: The teacher is a full-modality model with RGB and audio modalities. The student uses only the RGB modality, without KD, serving as the lower-bound baseline.
LogitKD [thoker2019cross, radevski2023multimodal]: Logit-level KD methods, in which the student learns from the teacher’s softened output logits using Kullback–Leibler divergence [hinton2015distilling].
ModalKD [liu2021semantics]: Feature-level KD method, in which the teacher network distills knowledge to the visual and pseudo-audio features of the student network.
DualKD [garcia2018modality]: Feature-level and logit-level KD method, in which the teacher network distills knowledge to features and logits of the student network.
CoKD (ViCoKD w/o Vi): The proposed ViCoKD method without the View-aware Consistency module.
Models & Hyperparameters. We adopt the backbone architecture and hyperparameters of the teacher and student networks from the original work. Following Nguyen et al. [nguyen2025multisensor], we sample input videos at 2.5 FPS and fix the sequence length to frames, based on the average video duration in the dataset. The cross-modal adapter is implemented as a lightweight feed-forward module with two linear layers and a Rectified Linear Unit (ReLU) [agarap2018deep] activation in between. The human-detection masks are generated by a You Only Look Once v10 (YOLO)-based [wang2024yolov10] detector. All loss terms are equally weighted with coefficients set to for simplicity. Optimization is performed using the Adaptive moment estimation (Adam) [kingma2014adam] optimizer with an initial learning rate of , weight decay of , and a batch size of for training epochs. The learning rate is scheduled using cosine annealing [loshchilov2016sgdr] for gradual decay over the course of training.
4.2 Quantitative Results
Table 1 presents the mAP results of the proposed ViCoKD method compared with competitive KD baselines across three backbones: MultiTrans [yasuda2022multi], MultiASL [nguyen2024action], and MultiTSF [nguyen2025multisensor]. The evaluation covers KD supervision scenarios: AV V and AV V, which denote distillation from an audio-visual teacher to a visual-only student with frame-level and sequence-level supervisions, respectively.
Home-1 Environment. This environment is particularly challenging due to its greater view disparity and weaker inter-view correlation. ViCoKD consistently delivered the highest mAP across all backbones and both KD supervision levels, with particularly large gains in the sequence-level regime. Using the MultiASL [nguyen2024action] backbone, ViCoKD improved mAP over the non-distilled student by (AV V) and (AV V), surpassing the teacher in both cases. The gains over CoKD (ViCoKD w/o Vi) across all experiments for both frame-level and sequence-level supervisions highlight the contribution of the proposed View-aware Consistency module in leveraging spatially localized cues under partial overlaps. In contrast, methods such as ModalKD and DualKD often underperformed the teacher, while LogitKD occasionally performed worse than the non-distilled student, underscoring their inability to adapt to scenarios where cross-view correspondence is weak or sparse.
Home-2 Environment. In this environment, the sensor views are closer in what they capture, which makes the setting less challenging than Home-1. Under these conditions, ViCoKD still secured top performance across all backbones. It achieved mAP improvements of up to (AV V) and (AV V) over the non-distilled student baseline. While LogitKD, ModalKD, and DualKD performed competitively in AV V, their advantages diminished under sequence-level supervision (AV V), reflecting the difficulty of exploiting weak labels. The consistent margins over CoKD reaffirm the effectiveness of the View-aware Consistency module in integrating human-centric spatial constraints into partially overlapping setups.
Home-1 | Home-2 | |||
| AV V | AV V | AV V | AV V | ||
| MultiTrans | 66.27 0.59 | 62.66 0.77 | 90.86 0.36 | 89.31 0.40 | |
| MultiASL | 76.83 0.80 | 65.23 0.74 | 91.80 0.85 | 89.89 0.57 | |
| MultiTSF | 82.91 0.65 | 65.28 0.21 | 91.27 0.79 | 89.21 0.70 | |
Key Findings. Across environments and backbones, ViCoKD consistently attained the highest performance, with particularly large gains in sequence-level supervision (AV V), where label sparsity is most severe. Moreover, as shown in Table 2, ViCoKD achieved high mAP while maintaining low variance ( of mAP) over three runs, indicating stable convergence and robustness. Figure 3 further demonstrates that ViCoKD converged faster and reached higher mAP than LogitKD and CoKD across all supervision settings and environments. The clearest margin was observed under AV V in both environments, where ViCoKD surpassed the non-distilled student early in training and maintained high performance, while LogitKD and CoKD achieved lower scores or exhibit fluctuations. In addition, Table 3 shows that applying the View-aware Consistency module to the teacher network itself yielded consistent mAP improvements. These results indicate the general effectiveness of the module as a principled mechanism for improving multi-view representation learning.
Home-1 | Home-2 | |
| MultiTrans [yasuda2022multi] | 61.40 66.31 | 86.60 88.88 | |
| MultiASL [nguyen2024action] | 73.81 75.99 | 90.14 91.78 | |
| MultiTSF [nguyen2025multisensor] | 76.12 82.16 | 92.12 92.61 |
4.3 Ablation Studies
We conducted ablation studies to quantify the contribution of each component in the proposed ViCoKD method, with the experimental results shown in Table 4. Note that some changes resulted in differences smaller than of mAP, which can be considered minor and likely within run-to-run variation, while others caused substantial drops, as indicated in Table 2.
View-aware Consistency Module. Eliminating Confidence Weighting (w/o ConfW) or the human-detection mask (w/o Mask) yielded notable drops of up to of mAP, particularly in the Home-1 environment within AV V. This indicates that weighting supervision by prediction reliability and masking irrelevant views are both crucial for robust cross-view alignment. Replacing the divergence from JS to KL (JS KL) also caused a small performance drop, showing the benefit of using a symmetric divergence.
Knowledge Distillation Strategies. Removing Feature-level Distillation (w/o FD) resulted in small drops in most settings, whereas omitting Logit-level Distillation (w/o LD) caused a large decrease in AV V. This suggests that while both FD and LD contributed to the overall improvement, LD was particularly important in scenarios with only sequence-level supervision.
Home-1 | Home-2 | |||
| AV V | AV V | AV V | AV V | ||
| ViCoKD | 76.83 | 65.23 | 91.80 | 89.89 | |
| View-aware Consistency Module | |||||
| w/o ConfW | 73.15 | 62.27 | 90.51 | 88.90 | |
| w/o Mask | 72.33 | 63.34 | 90.08 | 89.19 | |
| JS KL | 75.50 | 64.87 | 90.11 | 89.62 | |
| Knowledge Distillation Strategies | |||||
| w/o FD | 76.52 | 64.85 | 90.44 | 88.90 | |
| w/o LD | 76.23 | 61.51 | 92.02 | 88.76 | |
4.4 Qualitative Results
Figure 4 compares the attention maps of the teacher, the baseline non-distilled student, and the proposed ViCoKD for two representative actions across multiple sensor views. In the Home-1 environment example (“enter the room”), the baseline student often exhibited diffuse or background-focused attention, missing the key human regions. In contrast, ViCoKD consistently produced sharp, human-centric activation that tightly aligned with the actor’s location across all views, even when the viewpoint changed substantially. Similarly, in the Home-2 environment example (“stand up”), the attention of the student was scattered and shifted toward irrelevant areas such as furnitures. In contrast, ViCoKD focused precisely on the subject’s body, capturing the motion cues critical for action recognition. In several views, ViCoKD attended more precisely to human regions where the action occured, indicating that the proposed method transfered and refined the teacher’s learned features, enabling the model to focus on discriminative regions despite partial view overlaps and reduced modalities.
5 Limitations and Future Work
While the proposed ViCoKD method demonstrated strong performance in partially overlapping multi-view action recognition, it has several limitations. First, the framework relies on pre-trained human detectors to generate human-detection masks, which may introduce errors under heavy occlusion, poor lighting, or unconventional poses. These errors can affect the reliability of View-aware Consistency supervision. Second, the method depends on a strong multi-modal teacher, which limits applicability in domains where high-quality multi-modal data are scarce or unavailable. Finally, the current experiments are limited by available multi-view action recognition datasets, which typically provide only audio and visual modalities and lack other informative sources such as depth or skeleton signals.
Future work will explore end-to-end learned visibility estimation for improved robustness, as well as extending ViCoKD to additional modalities and cross-domain generalization.
6 Conclusion
We addressed the challenge of multi-view action recognition in partially overlapping sensor setups under modality and annotation-limited conditions. We introduced View-aware Cross-modal Knowledge Distillation (ViCoKD), a novel knowledge distillation method that transfers supervision from a fully supervised multi-modal teacher to a constrained student through cross-modal attention and a view-aware consistency mechanism. The proposed human-detection masks with confidence-weighted Jensen–Shannon divergence [menendez1997jensen] ensure that distillation focuses on frames with reliable, view-consistent evidence. Extensive experiments on the MultiSensor-Home [nguyen2025multisensor] dataset demonstrated that ViCoKD achieved substantial gains over non-distilled students and competitive baselines, with consistent improvements across backbones and environments. These results highlight the importance of explicitly modeling View-aware Consistency in partially overlapping scenarios and provide a foundation for more robust multi-view action recognition in real-world deployments.
Acknowledgment
This work was partly supported by JSPS KAKENHI JP21H03519 and JP24H00733. The computation was carried out using the General Projects on supercomputer “Flow” at IT Center, Nagoya University.
Appendix A Experimental Conditions
A.1 Datasets
We conduct the experiments on the MultiSensor-Home dataset [nguyen2025multisensor], a large-scale benchmark for multi-modal and multi-view human action recognition in real-world home environments. As shown in Figure 5, each home environment is equipped with multiple synchronized RGB and audio sensors positioned to capture daily activities from different viewpoints with partial overlaps. This multi-modal multi-view setup enables both frame-level and sequence-level supervision, but also introduces challenges due to varying degrees of inter-view alignment and occlusion.
The dataset comprises two distinct home environments, denoted Home-1 and Home-2, whose statistics are summarized in Table 5. Video recordings were captured at a resolution of pixels and a frame rate of 30 FPS. The dataset consists of 5,250 untrimmed video clips, including 2,550 from Home-1 and 2,700 from Home-2, corresponding to 1,050 multi-view sequences. Each individual view contributes approximately 9.6 hours of footage in Home-1 and 11.2 hours in Home-2. With five synchronized sensors deployed in each environment, the dataset totals approximately 20.8 hours of multi-view recordings. The average recording length is 68 seconds in Home-1 and 74 seconds in Home-2. Both environments have 16 action categories (Table 6), covering a wide range of common household activities such as “enter the room”, “sit down”, “use laptop”, and “open curtain”. Some classes are environment-specific, for example, “AdjustAC” appears only in Home-1, while “WatchTV” is unique to Home-2.
| Attribute | Home-1 | Home-2 |
| Total # of Multi-view Sequences | 510 | 540 |
| Total # of Videos | 2,550 | 2,700 |
| Total Multi-view Sequence Duration | 9.6 hours | 11.2 hours |
| Total Duration | 48 hours | 56 hours |
| Average Video Duration | 68 seconds | 74 seconds |
| Average # of Actions in Video | 3 | 3 |
| # of Action Classes | 16 | 16 |
| Total # of Actions | 1,334 | 1,171 |
The two environments differ in spatial configuration and view correlation: Home-1 exhibits greater viewpoint disparity and weaker cross-view correspondence, making it more challenging for cross-view knowledge transfer, whereas Home-2 offers stronger inter-view alignment. These characteristics make MultiSensor-Home a suitable testbed for evaluating the robustness of knowledge distillation under varying degrees of view overlap and label sparsity.
| Classes | Description | #Home-1 | #Home-2 |
| AdjustAC | Adjusting air conditioner | 39 | — |
| Clean | General cleaning activity | 26 | 33 |
| CleanVacuum | Cleaning with vacuum cleaner | 48 | 45 |
| OpenCurtain | Opening the curtain | 38 | 35 |
| CloseCurtain | Closing the curtain | 39 | 35 |
| Drink | Drinking water | 51 | 59 |
| Eat | Eating food | 48 | 51 |
| Enter | Entering the room | 70 | 82 |
| Exit | Exiting the room | 88 | 82 |
| ReadBook | Reading a book | 64 | 79 |
| Sitdown | Sitting down | 247 | 142 |
| Standup | Standing up | 161 | 126 |
| TurnOnLamp | Turning on the lamp | 57 | 49 |
| TurnOffLamp | Turning off the lamp | 52 | 41 |
| UseLaptop | Using a laptop computer | 196 | 114 |
| UsePhone | Using a phone | 110 | 102 |
| WatchTV | Watching the television | — | 96 |
| Total | 1,334 | 1,171 | |
A.2 Evaluation Metrics
Following [kobayashi2023two, nguyen2024action, nguyen2025multisensor], we evaluate performance using the mean Average Precision (mAP) based on a macro-averaged metric, which is widely used in multi-label classification tasks. mAP is computed independently for each class over samples and then aggregated across all classes as:
where AP denotes the Average Precision computed for a set of predictions and ground truths for class . This metric effectively evaluates how well the model ranks relevant instances ahead of irrelevant ones across multiple labels.
A.3 Resource
All experiments are conducted on a FUJITSU CX2570M5 server equipped with two Intel Xeon Gold 6230 CPUs and four NVIDIA Tesla V100 GPUs.
Appendix B Experimental Results
In this section, we provide additional experimental results to further evaluate the effectiveness of the proposed Cross-modal Knowledge Distillation (ViCoKD) method.
B.1 Testing mAP Curves under Different Distillation Settings
Figures 6 and 7 present the testing mAP curves for the MultiTrans [yasuda2022multi] and MultiTSF [nguyen2025multisensor] backbones under various distillation strategies. A consistent trend emerges across both architectures: the proposed ViCoKD method converged more rapidly and attained higher or more stable mAP compared to LogitKD and CoKD. In particular, under the AV V setting, ViCoKD demonstrated clear advantages by surpassing the non-distilled student early in training and sustaining strong performance throughout.
For MultiTrans [yasuda2022multi] backbone (Figure 6), the improvements were most noticeable in the Home-1 environment, where other methods showed slower convergence or larger fluctuations. For MultiTSF [nguyen2025multisensor] backbone (Figure 7), ViCoKD not only delivered the best final performance but also reduced variance across epochs, suggesting more stable optimization dynamics.
Overall, these additional results confirm that the effectiveness of ViCoKD is not limited to a single backbone but generalized well across different temporal fusion architectures. This reinforces the role of View-aware Consistency in guiding robust representation learning under partial cross-modal supervision.
B.2 Qualitative Results
In addition to the qualitative results presented in the main text, Figure 8 provides further visualizations of attention maps for diverse actions. Across both Home-1 and Home-2 environments, the proposed ViCoKD method consistently produced human and the action-related attention than either the baseline student or the teacher.
Home-1 environment: For the “sit down” action (Figure 8(a)) across Views 1, 2, and 3, the teacher model distributed attention over the actor’s body posture changes, while the baseline student often drifted toward background elements such as furniture or walls. ViCoKD, in contrast, narrowed its focus directly on the seated human across all views, maintaining cross-view consistency. Similarly, in the “clean by vacuum” action (Figure 8(c)) on Views 4 and 5, the teacher highlighted the person in View 4 and the vacuum cleaner in View 5. ViCoKD recovered this alignment by highlighting both the human and the vacuum cleaner in both views, reliably centering on the human and the action-related tool across perspectives.
Home-2 environment: For the “watch TV” action (Figure 8(b)), the teacher mainly attended to the TV screen but often neglected the human actor, especially in View 5. The baseline student, in contrast, produced scattered and weak attention maps, frequently drifting toward irrelevant floor regions. ViCoKD recovered a more balanced focus, consistently attending to both the TV and the human subject across all views, thereby capturing the essential human-object interaction that defines the action. For the “use laptop” action (Figure 8(d)), the attention maps of the teacher, baseline student, and ViCoKD appear more similar compared to other actions. All three models highlighted the subject and the laptop region to some degree, reflecting the strong visual saliency of the interaction.